### v1

@@ 3 Acoustic detection
Figure 3: The critical detectable distance $R_{\rm crit}$. For different $A$ and $k$, the critical detectable distance is similar, about 30 km shown by the solid line.
@@ 4.1 Distinction of SN signals
$x_{\rm pene}=110\,{\rm cm}\cdot(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm{dust}}}{3\,\rm{g\cdot cm^{-3}}})^{2/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1},$ (18)
@@ 4.2 Detection efficiency of the acoustic energy
As the SN passes through and compresses the medium, some of the kinetic energy deposited is used to generate a sound wave, and some is used to heat and ionise the medium. In addition, the sound wave will reflect on the surface of microphone box or the hydrophone. We can define a factor $\eta$ to describe the detection efficiency of the acoustic energy, $\mathrm{d}E_{\rm s}\rightarrow\eta\mathrm{d}E_{\rm s}$, the Equation (17) becomes
$\begin{split}\exp(-\dfrac{R_{\rm crit}}{1\,\mathrm{km}})&=\dfrac{4\pi R_{\rm crit}p_{\rm noise}^{2}t_{\rm s}}{\eta\rho_{\mathrm{med}}^{2}c_{\rm s}\sigma_{\rm SN}v^{2}}\\ &=\dfrac{1}{\eta}\exp(-19.88)\cdot(\dfrac{A}{10^{30}})^{-5/12}(\dfrac{R_{\rm crit}}{1\,\mathrm{km}})^{5/4}(\dfrac{p_{\rm noise}}{1\,\rm{erg\cdot cm^{-3}}})^{2}(\dfrac{v}{220\,\rm{km\cdot s^{-1}}})^{-7/4}\\ &\cdot k^{-5/8}(\dfrac{c_{\rm s}}{1\,\rm{km\cdot s^{-1}}})^{-9/4}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-2}.\end{split}$ (19)
However, the influence of $\eta$ on the critical detectable distance $R_{\rm crit}$ is small. Figure 4 shows the critical detectable distance when $\eta=10^{-6}$, it is about half of the $R_{\rm crit}$ when $\eta=1$.
Figure 4: The critical detectable distance $R_{\rm crit}$, $\eta=10^{-6}$. It is about 1/3 to 1/2 of $R_{\rm crit}$ when $\eta=1$.
@@ 4.3 The accumulation of SNs in the Earth or the Moon
We can place an upper bound on the accumulation rate and the total accumulated mass of SNs during the Earth’s 4.6 Gyr lifetime. From Figure 1, SN with about $A\leq 10^{30},k\geq 1$ could not pass through the Earth and may sink to the center of the Earth. From Equation (2), we can calculate that the accumulation rate is less than $4.8\times 10^{8}\,{\rm g\cdot yr^{-1}}$, the total accumulated mass is $2.2\times 10^{18}\,{\rm g}\simeq 3.7\times 10^{-10}M_{\oplus}$. The radius of the “SN core” in the Earth is $17$ cm. The gravitational effect is tiny. The accumulation rate in the Moon is about 1/50 that of the Earth.
@@ 4.5 Other uses of the acoustic array
For earthquakes, the existing seismic stations operate in the frequency band less than or equal to 1 kHz (Tian & Liang 2013; Goldman et al. 2020; Chen et al. 2022; Darzi et al. 2024). The acoustic array can extend the seismic detection band up to 96 kHz. The distributed acoustic sensing (DAS) technology can combine acoustic and seismic detection (Parker et al. 2014; Ajo-Franklin et al. 2019; Xenaki et al. 2025). It can be used to acquire the Earth’s motion field over a certain area, which is related to SN impact events and earthquakes. Some studies have focused on approaches to placing acoustic sensors so that they can acquire data more efficiently (Saad et al. 2024; Callahan et al. 2024).
@@ 4.6 The budget estimation
[…]
The material cost is discussed next. In order to reduce costs, we can use civilian 16-bit 192 kHz microphones. These will pick up enough sound to detect a 74 dB SN signal. The 192 kHz sampling rate is sufficient to detect an SN with $A\geq 10^{26}$ and $k\geq 0.1$ at a 30 km distance. One microphone is about RMB 50, one meter of cable is about RMB 1, one 5 µs time synchronisation module is about RMB 100, one wireless transmission module is about RMB 50, one directional antenna is about RMB 50, one hydrophone is about RMB 200, one buoy is about RMB 50. The data collection system and data processing computer may cost RMB 100 k. Based on the above description, the material cost of one site is RMB 560 to RMB 650 on the ground, or RMB 650 in the ocean. So the material cost of the array is about RMB 155 k to RMB 165 k.
@@
It is a mercy that the High-energy Underwater Neutrino Telescope (HUNT)  [^5] http://hunt.ihep.ac.cn/ has been proposed, which will use sonar to localize each underwater neutrino detector. The HUNT will contain about 2000 vertical detector strings at 150 m horizontal intervals, each string will include at least 4 sonar localisation modules. The array design of HUNT is $6\,{\rm km}\times 6\,{\rm km}\times 1\sim 2\,{\rm km}$ at a depth of greater than 1 km. The acoustic localisation system on the HUNT can also detect the SN signal with an event rate of about $1\,{\rm yr^{-1}}\cdot(A/10^{26})^{-1}$. The acoustic data from HUNT would certainly constrain the baryon number $A$ of SNs in the future.

### v2

@@ 3 Acoustic detection
[…]
@@ 4.1 Distinction of SN signals
$\begin{split}x_{\rm pene}&=110\,{\rm cm}\\ &\cdot(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm{dust}}}{3\,\rm{g\cdot cm^{-3}}})^{2/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1},\end{split}$ (11)
@@ 4.2 Detection efficiency of the acoustic energy
As the SN passes through and compresses the medium, some of the kinetic energy deposited is used to generate a sound wave, and some is used to heat and ionise the medium. In addition, the sound wave will reflect on the surface of the microphone box or the hydrophone. We can define a factor $\eta$ to describe the energy transfer efficiency from kinetic energy lost by the SN to detected acoustic energy, the energy transfer relation $\mathrm{d}E_{\rm s}=-\mathrm{d}E_{\rm k}$ is then changed to $\mathrm{d}E_{\rm s}=-\eta\cdot\mathrm{d}E_{\rm k}$, the Equation (10) becomes
$\begin{split}\exp(-\dfrac{R_{\rm crit}}{1\,\mathrm{km}})&=\dfrac{4\pi R_{\rm crit}p_{\rm noise}^{2}t_{\rm s}}{\eta\rho_{\mathrm{med}}^{2}c_{\rm s}\sigma_{\rm SN}v^{2}}\\ &=\dfrac{1}{\eta}\exp(-16.03)(\dfrac{A}{10^{30}})^{-5/12}(\dfrac{R_{\rm crit}}{1\,\mathrm{km}})^{5/4}\\ &\cdot(\dfrac{p_{\rm noise}}{1\,\rm{erg\cdot cm^{-3}}})^{2}(\dfrac{v}{220\,\rm{km\cdot s^{-1}}})^{-7/4}\\ &\cdot k^{-5/8}(\dfrac{c_{\rm s}}{1\,\rm{km\cdot s^{-1}}})^{-9/4}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-2}.\end{split}$ (12)
However, the influence of $\eta$ on the critical detectable distance $R_{\rm crit}$ is small. Figure 4 shows the critical detectable distance when $\eta=10^{-6}$, it is about half of the $R_{\rm crit}$ when $\eta=1$. Mathematically speaking, the change of $\eta$ is equivalent to the change of $p_{\rm noise}$, $v$, $c_{\rm s}$, or $\rho_{\rm med}$, so the different values of them will not have a significant impact on the $R_{\rm crit}$.
Figure 4: The critical detectable distance $R_{\rm crit}$ when the energy transfer efficiency $\eta$ from kinetic energy lost by the SN to detected acoustic energy is $10^{-6}$. It is about 1/3 to 1/2 of $R_{\rm crit}$ when $\eta=1$.
@@ 4.3 The accumulation of SNs in the Earth or the Moon
We can place an upper bound on the accumulation rate and the total accumulated mass of SNs during the Earth’s 4.6 Gyr lifetime. From Figure 1, SN with about $A\leq 10^{30}$ could not pass through the Earth and may sink to the center of the Earth. From Equation (2), we can calculate that the accumulation rate is less than $4.8\times 10^{8}\,{\rm g\cdot yr^{-1}}$, the total accumulated mass is $2.2\times 10^{18}\,{\rm g}\simeq 3.7\times 10^{-10}M_{\oplus}$. The radius of the “SN core” in the Earth is $17$ cm. The gravitational effect is tiny. The accumulation rate in the Moon is about 1/50 that of the Earth.
@@ 4.5 Other uses of the acoustic array
For earthquakes, the existing seismic stations operate in the frequency band less than or equal to 1 kHz (Tian & Liang 2013; Goldman et al. 2020; Chen et al. 2022; Darzi et al. 2024). The acoustic array can extend the seismic detection band up to 96 kHz. The distributed acoustic sensing (DAS) technology can combine acoustic and seismic detection (Parker et al. 2014; Ajo-Franklin et al. 2019; Xenaki et al. 2025). It can be used to acquire the Earth’s motion field over a certain area, which is related to SN impact events and earthquakes. Some studies have focused on approaches to placing acoustic sensors so that they can acquire data more efficiently (Saad et al. 2024; Callahan et al. 2025).
@@ 4.6 The budget estimation
First we need to clarify the composition of the array. Building a single $670\,{\rm km}\times 670\,{\rm km}$ microphone array is unattainable, but we can build a small array that will be used first for technology validation and preliminary detection. If we build a microphone array horizontally to $100\,{\rm km}\times 100\,{\rm km}$, using a 10-kilometer square grid, we need $100$ acoustic observatory sites. On the ground, each site contains a microphone box with 5 microphones at a depth of 10 m to 100 m in the bedrock. In the ocean, each site contains a hydrophone at a depth of 200 m, and a buoy to place the electronics equipment and antenna. Each site includes a wireless transmission module that encodes and gives data to the antenna, a directional antenna that transmits data, and a synchronised clock accurate to 5 µs to ensure that the signals from different microphones are synchronised. We also need a data collection system for the whole array. The maximum data rate is $192\,{\rm kHz}\times 16\,{\rm bits\ per\ pressure\ sample}\times 100\times 2=614.4\,{\rm Mbit\cdot s}^{-1}$, it is achievable with a 1 Gbps data broadband.
The material cost is discussed next. In order to reduce costs, we can use civilian 16-bit 192 kHz microphones. These will pick up enough sound to detect a 74 dB SN signal. The 192 kHz sampling rate is sufficient to detect an SN with $A\geq 10^{26}$, $k\geq 10$ at a 30 km distance. One microphone is about RMB 50, one meter of cable is about RMB 1, one 5 µs time synchronisation module is about RMB 100, one wireless transmission module is about RMB 50, one directional antenna is about RMB 50, one hydrophone is about RMB 200, one buoy is about RMB 50. The data collection system and data processing computer may cost RMB 100 k. Based on the above description, the material cost of one site is RMB 560 to RMB 650 on the ground, or RMB 650 in the ocean. So the material cost of the array is about RMB 155 k to RMB 165 k.
@@
It is a mercy that the High-energy Underwater Neutrino Telescope (HUNT)  [^4] http://hunt.ihep.ac.cn/ has been proposed, which will use sonar to localize each underwater neutrino detector. The HUNT will contain about 2000 vertical detector strings at 150 m horizontal intervals, each string will include at least 4 sonar localisation modules. The array design of HUNT is $6\,{\rm km}\times 6\,{\rm km}\times(1\sim 2)\,{\rm km}$ at a depth of greater than 1 km. The acoustic localisation system on the HUNT can also detect the SN signal with an event rate of about $1\,{\rm yr^{-1}}\cdot(A/10^{26})^{-1}$. The acoustic data from HUNT would certainly constrain the baryon number $A$ of SNs in the future.
@@ Acknowledgements.
Appendix A The cross section of different kinds of SN
We are concerned about the cross section of SN interacting with ordinary matter below. As stated in Equation (4), the geometric cross section of an SN is
$\begin{split}\sigma_{\mathrm{SN\_geo}}&=8.0\times 10^{-6}\,\rm{cm}^{2}\\ &\cdot(\dfrac{\it{A}}{10^{30}})^{2/3}(\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{-2/3}.\end{split}$
[…]
where $\alpha_{\rm em}\equiv e^{2}/(\hbar c)\simeq 1/137$. As $\Delta p\ll p_{\rm F}$, there is only a small fraction of electrons whose spins are parallel when the Coulomb interaction is included for an electron gas. Therefore, the total magnetic moment reads,
where $\mu_{\rm B}=e\hbar/(2m_{\rm e})$ is the Bohr magneton and $N_{\rm e}$ is the total number of electrons in an SnN. The magnetic field on the magnetic equator of the SnN is thus,
$B_{0}=\dfrac{\xi\mu}{r_{0}^{3}}=8.4\times 10^{11}\,\rm{G}\cdot\xi\dfrac{\delta_{\rm e}}{10^{-5}}\cdot\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}},$ (17)
which is close to the typical magnetic fields on pulsars’ surface [^6] This mechanism of spontaneous magnetisation is different from that of conventional ferromagnetism. The ideal fermion gas approximation is perfect for the former, extremely relativistic electrons, but not for the latter, non-relativistic electrons in metals. Complex interactions matter in metals. .
When an SN impacts the ordinary matter, such as soil, water or rock, it makes a deposition of its kinetic energy into the surrounding medium. The way SN loses energy is by compressing the medium and transferring momentum to the molecules of the medium. Some of the energy is converted to heat, creating an ionised environment, and some of the energy creates the shock wave as an acoustic signal. In the aerodynamic heating picture, the temperature rises, $\Delta T=v^{2}/2c_{\rm p}$, where $c_{\rm p}$ is the specific isobaric heat capacity. For solid SiO2, $c_{\rm p}=7.0\times 10^{6}\,{\rm erg\cdot(g\cdot K)^{-1}}$. Then substitute $v=220\,\rm{km\cdot s^{-1}}$, we have $\Delta T=7.0\times 10^{7}\,$, which is high enough to ionise the atoms, causing the medium surrounding the impact trajectory to become plasma. It is worth mentioning that SNs will not absorb ordinary matter into itself. Firstly, the “nuclear charge number” of an SN is $Z=\delta_{\rm e}A\simeq 10^{20}\sim 10^{30}$. Even when considering the nuclear charge shielding caused by the Thomas-Fermi distribution of electrons, the Coulomb potential energy barrier is too high for non-relativistic ordinary atomic nuclei to overcome because the geometric radius $r_{0}$ of the SN is far greater than the electron Compton wavelength $\lambda_{\rm C}$. Secondly, for nuggets with strangeness, ordinary matter must to be transformed into strange matter via weak interactions before it can be absorbed. The probability of such a transformation occurring in a single collision is extremely low.
$\begin{split}r_{\rm m}&=(\dfrac{B_{0}^{2}r_{0}^{6}}{4\pi\rho_{\rm med}v^{2}})^{1/6}\\ &=0.035\,{\rm cm}\cdot(\dfrac{\it{A}}{\mathrm{10^{30}}}\xi\dfrac{\delta_{\rm e}}{\rm{10^{-5}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1/6}\\ &\cdot(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-1/3}.\end{split}$ (18)
It implies that $r_{\rm m}\gg r_{0}$, and $r_{\rm m}/r_{0}=22\cdot(\dfrac{\xi\delta_{\rm e}}{\rm{10^{-5}}}\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{\frac{1}{3}}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{\frac{1}{6}}(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-\frac{1}{3}}$ is independent of the baryon number $A$, i.e. the size of SnN. Thanks to the large $r_{\rm m}$, acoustic detection of the SN is more feasible.
The cross section of the magnetosphere of the SnN reads,
$\begin{split}\sigma_{\rm m}&=\pi r_{\rm m}^{2}\\ &=3.8\times 10^{-3}\,{\rm cm}^{2}\cdot(\dfrac{\it{A}}{\rm{10^{30}}}\xi\dfrac{\delta_{\rm e}}{\rm{10^{-5}}})^{2/3}\\ &\cdot(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1/3}(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-2/3},\end{split}$ (19)
which could be regarded as the maximum effective cross section of the SnN in the ordinary matter medium. Refer to Equation (5)
$\sigma_{\rm SN}=k\cdot{(\dfrac{\it{A}}{\rm{10^{30}}})}^{2/3}\cdot 8.0\times 10^{-6}\,{\rm cm}^{2},$
Equation (19) is equivalent to that $k=4.8\times 10^{2}$ for SnNs. Considering that parameters $\delta_{\rm e}$ and $\rho_{\rm med}$ may vary, and the magnetosphere may not completely block particles from passing through, $k=10^{1}\sim 10^{2}$.
As with the SnN, the charge-mass ratio of a strange quark nugget is also $\delta_{\rm e}=10^{-6}\sim 10^{-5}$. Its electrons can also magnetize spontaneously, so $k=10^{1}\sim 10^{2}$ for strange quark nugget. The acoustic signals of SnNs and strange quark nuggets are indistinguishable. However, in the case of the $ud$-quark nugget, flavor asymmetry leads to a higher charge-mass ratio. Using the MIT bag model, we can calculate that $\delta_{\rm e}$ can be up to $10^{-2}$ for the $ud$-quark matter, which will result in a greater dipole magnetic field and a larger radius of the magnetosphere. So $k\simeq 10^{3}$ for the $ud$-quark nugget. For the axion-quark nuggets, since the flavor of quarks they contain is unknown and the effects of axions are unknown, their cross sections are uncertain. We take $k=10^{-1}\sim 10^{3}$, from incompletely momentum transfer via the geometric cross section case to the largest magnetosphere case.